• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 23:04:05 on October12,2005

    Entry number 156677

    keyword=spectr. DAQ runs

    Beam 0.2 uA (actually see 0.28 on EPICS).
    LH2 target in.
    Switched to spectrometer DAQ. Did the checklist.
    Rates:
    T1 = 195 kHz
    T2 = 145 kHz
    T3 = 192 kHz
    T4 = 128 kHz
    T5 = T6 = 0
    T7 = 259 kHz
    T8 = 243 kHz

    Spectrometer run 2814.
    P_L = 3.128, P_R = 3.127 nominally correct.
    All HAPPEX detector 1900 V.

    Problem: One of the HAPPEX detectors is missing. It appears
    in neither ADC nor TDC nor scaler.
    The detector map db_P.dat says it should be in the
    3rd descriminator input, which is consistent with halog 156639
    but it looks like it is not plugged in.

    Meanwhile we put momentum to -2% to start checking alignment but
    we'll need to take access to fix the above, before going too far.





    FIGURE 1